A sample of blood pressure measurements is taken for a group of​ adults, and those values​ (mm Hg) are listed below. The values are matched so that 10 subjects each have a systolic and diastolic measurement. Find the coefficient of variation for each of the two​ samples; then compare the variation. 119,126,157,96,154,123,115,135,125,118


1
Expert's answer
2021-09-27T03:05:08-0400

"\bar{x} = \frac{\sum x}{n} \\

\bar{x} = \frac{119+126+157+96+154+123+115+135+125+118}{10}=127.1 \\

s = \sqrt{\frac{\sum(x-\bar{x})^2}{n-1}} \\

s = \sqrt{\frac{(119-127.1)^2 + (129-127.1)^2+...+(118-127.1)^2}{9}} \\

= \frac{2966.9}{9} = 329.65"

Coefficient of variation "= \frac{s}{\bar{x}} = \frac{127.1}{329.65} = 0.385"